# Defending Large Language Models Against Attacks With Residual Stream Activation Analysis thanks: Research sponsored by HiddenLayer

### v3

@@ Abstract
Index Terms:
adversarial machine learning, large language models, machine learning
I Introduction
@@
In this paper, we refer to these malicious prompts as “attack prompts,” a broad term encompassing various forms of prompt-based intrusions including prompt injections, such as the attacks from Perez and Ribeiro, and jailbreaks, such as the “Developer Mode” attack [1],[3]. It is important to clarify that the term “attack prompts” as we define it does not refer to ”adversarial perturbations,” a style of attack associated with making minute, often imperceptible changes to input data in order to manipulate gradient calculations with the intent to force a machine learning model to misclassify a seemingly non-corrupted input [4].
@@
Section II describes the previous research that inspired this work. Section III describes our methodology for activation classification and Section IV details our LLM fine tuning methodology for increasing the accuracy of activation classification. Section V discusses our results and final remarks are presented in Section VI.
II Related Work
@@
Templeton et al (2024) trained autoencoders on the Claude 3 LLM activation space in order to discover features of abstract concepts, including concepts such as ”security vulnerabilities in code” and ”bias” [8]. We use a simpler model to classify a spectrum of attack behaviors as specified by a potential user. This specification is provided through a example prompts, rather than abstract concepts.
III Methodology - Activation Classification
In this section, we define residual stream activations in a transformer model (III-A) and detail the steps we take to analyze and classify them (III-B). We also describe the malicious and benign prompt datasets we use (III-C) and the LLMs we use in the experiments (III-D).
InputTransformer Layer 1Residual Connection 1Transformer Layer 2Residual Connection 2Transformer Layer 3Residual Connection 3Residual Stream
@@ III Methodology - Activation Classification
III-A Residual Stream Activations
@@
III-B Activation Analysis Methodology
@@
    Activation Vector Collection: For a prompt $P_{m}$ with $N$ tokens $t_{1},t_{2},\ldots,t_{N}$, where $m$ indexes the prompt from 1 to $M$ (the total number of prompts in a dataset), we collect activation vectors from each layer. For each token $t_{k}$ in prompt $P_{m}$, we obtain 32 activation vectors (one per layer), denoted as $v_{1}(t_{k}),v_{2}(t_{k}),\ldots,v_{32}(t_{k})$.
@@
    $v_{1}^{(m)}=\text{average}(v_{1}(t_{1}),v_{1}(t_{2}),\ldots,v_{1}(t_{N})),$
@@
    $v_{32}^{(m)}=\text{average}(v_{32}(t_{1}),v_{32}(t_{2}),\ldots,v_{32}(t_{N})).$
    This averaging ensures that each set of activations for prompt $P_{m}$ has consistent dimensions regardless of the prompt’s length.
@@
    Classifier Training: We train a LightGBM classifier for each set of layer vectors to determine if the activations can be differentiated by class [5]. In this case, we train 32 classifiers, $\{C_{1},C_{2},\ldots,C_{32}\}$. Classifier $C_{i}$ uses set $V_{i}$ for training, where $V_{i}$ is the set of all $v_{i}^{(m)}$ vectors for every prompt $P_{m}$. The classifiers are configured with a maximum depth of 6 and 100 estimators. We reserve 10% of the activation dataset for testing to verify the generalizability of the classifiers and to check for overfitting.
TABLE I: Prompt Length Restriction by Dataset
@@
Domain-Specific Attack [10-845] [50-170] 4138
Benign [9-438] [50-170] 4138
Hyper-Specific Attack [63-606,391] [500-1500] 5358
Benign [22-35,871] [500-1500] 5358
III-C Data Sources
For our analysis, we use 3 different datasets, each encompassing a different scope of attack prompts. Our intention is to simulate attacks that make up a representative selection of the spectrum of possible attack domains. The different types of attacks are: Broad, Domain-Specific, and Hyper-Specific. The Broad category consists of attacks that attempt to engage an LLM in a broad spectrum of undesirable behaviors. The Domain-Specific category consists of attacks pertaining to a single domain, simulating attacks on an LLM that is fine-tuned for a specific task. The Hyper-Specific category consists of attacks that aim at producing a specific undesirable string in an LLM’s response. Notably, we did not require that all attack prompts be successful. This is because an ideal LLM defense system should be able to detect on-going attacks on the model even if they are unsuccessful, so counter-measures can be taken. That said, we do test our classifier on an LLM that underwent fine-tuning to disengage safety training and on which all attacks were successful, further detailed in III-D.
III-C1 Broad Category
The attack dataset we use to encompass a broad range of attacks is the JailbreakV-28K dataset from HuggingFace [9]. We use the jailbreak_query column as the attacks. The dataset contains categories for each type of attack such as “Fraud”, “Animal Abuse”, and “Malware”, among others, which demonstrates the large variance of harmful topics for attack prompt generation. The class of benign prompts originates from Open-Orca on HuggingFace, a large collection of benign prompts for LLM training [10]. Although both of the original datasets have over 100,000 prompts each, we use a subset of the dataset for analysis due to time and computation constraints. The subset for our analysis contains 50,000 prompts split evenly between attack and benign prompts. We also reserve a smaller subset of the prompts for our fine tuning procedure: another 2,000 prompts split evenly between attack and benign prompts.
III-C2 Domain-Specific Category
TABLE II: Broad Attack Results - Averaged Over Layers
Model Accuracy Std Deviation Prompt Class Precision Recall
LLaMA 2 7B 0.99999 0.00003 Attack 0.99997 1.00000
Benign 1.00000 0.99997
LLaMA 2 13B 0.99973 0.00046 Attack 0.99971 0.99974
Benign 0.99974 0.99972
TinyLlama 1.1B 0.99941 0.00042 Attack 0.99920 0.99963
Benign 0.99963 0.99920
Mistral 7B 0.99999 0.00001 Attack 0.99999 1.00000
Benign 1.00000 0.99999
Vicuna 7B 0.99999 0.00002 Attack 0.99999 1.00000
Benign 1.00000 0.99999
Wizard Vicuna 7B Uncensored 0.99949 0.00051 Attack 0.99958 0.99939
Benign 0.99939 0.99958
TABLE III: Domain-Specific Attack Results - Averaged Over Layers
Model Accuracy Std Deviation Prompt Class Precision Recall
LLaMA 2 7B 0.99188 0.00555 Attack 0.99475 0.98898
Benign 0.98905 0.99478
LLaMA 2 13B 0.99445 0.00665 Attack 0.99785 0.99102
Benign 0.99110 0.99788
TinyLlama 1.1B 0.98602 0.00997 Attack 0.99048 0.98145
Benign 0.98167 0.99059
Mistral 7B 0.98663 0.00429 Attack 0.99469 0.97849
Benign 0.97886 0.99478
Vicuna 7B 0.98412 0.00771 Attack 0.99271 0.97537
Benign 0.97584 0.99286
Wizard Vicuna 7B Uncensored 0.99059 0.00858 Attack 0.99490 0.98621
Benign 0.98636 0.99496
TABLE IV: Hyper-Specific Attack Results - Averaged Over Layers
Model Accuracy Std Deviation Prompt Class Precision Recall
LLaMA 2 7B 0.99964 0.00035 Attack 0.99967 0.99961
Benign 0.99961 0.99967
LLaMA 2 13B 0.99956 0.00084 Attack 0.99973 0.99939
Benign 0.99939 0.99973
TinyLlama 1.1B 0.99938 0.00111 Attack 0.99942 0.99933
Benign 0.99933 0.99942
Mistral 7B 0.99977 0.00024 Attack 0.99974 0.99981
Benign 0.99981 0.99974
Vicuna 7B 0.99964 0.00054 Attack 0.99961 0.99967
Benign 0.99967 0.99961
Wizard Vicuna 7B Uncensored 0.99939 0.00084 Attack 0.99954 0.99924
Benign 0.99925 0.99954
III-C3 Hyper-Specific Category
The attack dataset for this analysis originates from submissions to HackAPrompt 2023, an online competition to create successful attack prompts [12]. The participants in this competition created prompts that attempted to make an LLM output the string:
@@ III-C3 Hyper-Specific Category
which is the goal output string that indicates that an LLM has been compromised. The attack dataset was then made open source on HuggingFace, a popular website for sharing models and datasets [13]. We use a subset of the original dataset, totaling 25,000 attacks. For the benign prompts, we use the same subset as the Broad dataset: 25,000 prompts from the Open-Orca dataset. We also use another smaller subset of the Hyper-Specifc data for our fine tuning procedure, another 2,000 prompts split evenly between attack and benign prompts.
We additionally perform prompt length range restriction on some of the datasets. The prompt classes in the Hyper-Specific and Broad datasets have different distributions of prompt length. In order to avoid prompt length becoming a determining feature of the classifier, we make modified versions of these datasets with the range of the prompt lengths restricted. The new versions of these datasets are subsets of the original datasets that only contain prompts that match the requirements for length. The new dataset is then class balanced by removing a random selection of prompts from the larger class. We perform our analysis with the original and range-restricted datasets and report results from each. Table I provides information on the transformations to each dataset. We determined that the prompt length range for the Domain-Specific dataset was similar enough between prompt classes and did not need any modification.
TABLE V: Broad Attack Results (Length Range Restricted) - Averaged Over Layers
Model Accuracy Std Deviation Prompt Class Precision Recall
LLaMA 2 7B 0.99951 0.00033 Attack 0.99943 0.99960
Benign 0.99960 0.99943
LLaMA 2 13B 0.99954 0.00061 Attack 0.99957 0.99950
Benign 0.99950 0.99957
TinyLlama 1.1B 0.99885 0.00066 Attack 0.99919 0.99851
Benign 0.99851 0.99919
Mistral 7B 0.99956 0.00030 Attack 0.99949 0.99962
Benign 0.99962 0.99949
Vicuna 7B 0.99956 0.00039 Attack 0.99942 0.99969
Benign 0.99969 0.99942
Wizard Vicuna 7B Uncensored 0.99953 0.00040 Attack 0.99935 0.99970
Benign 0.99970 0.99935
TABLE VI: Domain-Specific Attack Results (Length Range Restricted) - Averaged Over Layers
Model Accuracy Std Deviation Prompt Class Precision Recall
LLaMA 2 7B 0.99283 0.00510 Attack 0.99442 0.99124
Benign 0.99127 0.99442
LLaMA 2 13B 0.99510 0.00532 Attack 0.99689 0.99331
Benign 0.99333 0.99690
TinyLlama 1.1B 0.98904 0.01011 Attack 0.98893 0.98917
Benign 0.98917 0.98890
Mistral 7B 0.98929 0.00378 Attack 0.99184 0.98669
Benign 0.98677 0.99188
Vicuna 7B 0.99107 0.00559 Attack 0.99199 0.99014
Benign 0.99017 0.99200
Wizard Vicuna 7B Uncensored 0.99463 0.00625 Attack 0.99609 0.99316
Benign 0.99318 0.99609
TABLE VII: Hyper-Specific Attack Results (Length Range Restricted) - Averaged Over Layers
Model Accuracy Std Deviation Prompt Class Precision Recall
LLaMA 2 7B 0.99872 0.00124 Attack 0.99933 0.99811
Benign 0.99811 0.99933
LLaMA 2 13B 0.99870 0.00272 Attack 0.99912 0.99827
Benign 0.99827 0.99912
TinyLlama 1.1B 0.99858 0.00162 Attack 0.99788 0.99929
Benign 0.99929 0.99787
Mistral 7B 0.99930 0.00057 Attack 0.99963 0.99896
Benign 0.99896 0.99963
Vicuna 7B 0.99904 0.00115 Attack 0.99927 0.99881
Benign 0.99881 0.99927
Wizard Vicuna 7B Uncensored 0.99878 0.00150 Attack 0.99899 0.99857
Benign 0.99857 0.99899
III-D Models
@@
IV Methodology - Fine Tuning for Improved Classification
With the intention of improving attack prompt detection accuracy, we fine tune LLaMA 2 7B using Low Rank Approximation of LLMs (LoRA) [19]. This lightweight and parameter efficient training procedure allows us to fine tune the 7B model on one GPU. In total, we fine tune three versions of LLaMA 2 7B, once for each of the dataset types we detailed in III-C. For each model, we use a hold-out set of prompts solely used for fine tuning (2,000 for Broad, 1,000 for Domain-Specific, 2,000 for Hyper-Specific, each split evenly across classes). Taking inspiration from Raman and Maini (2023), we then apply a format to each prompt in order to force the model to evaluate each prompt in the context of detecting prompt injections [20]. We then provide responses to each prompt for the LLM to emulate. The LLM is then trained on the combination of augmented prompt and response to complete the LoRA fine tuning procedure. To make the augmented prompt, we started by prepending the string:
@@ IV Methodology - Fine Tuning for Improved Classification
to the end of the prompt. This frames the original prompt in an appropriate context and asks the LLM to only respond with “yes” or “no”. We then provide a corresponding response that the LLM should train to emulate. The response is “yes” if the original prompt belongs to the attack prompt class or “no” if the original prompt belongs to the benign prompt class. After fine tuning the three versions of LLaMA 2 7B, we repeat the activation classification procedure we detailed in III-B and compare the results to the results utilizing non-fine tuned LLMs.
V Results

### v4

# CAMLIS’24: Conference on Applied Machine Learning for Information Security, October 24–25, 2024, Arlington, VA

@@ Abstract
keywords:
[…]
1 Introduction
@@
In this paper, we refer to these malicious prompts as “attack prompts," a broad term encompassing various forms of prompt-based intrusions including prompt injections, such as the attacks from Perez and Ribeiro, and jailbreaks, such as the “Developer Mode” attack [1],[3]. It is important to clarify that the term “attack prompts" as we define it does not refer to ”adversarial perturbations," a style of attack associated with making minute, often imperceptible changes to input data in order to manipulate gradient calculations with the intent to force a machine learning model to misclassify a seemingly non-corrupted input [4].
@@
Section 2 describes the previous research that inspired this work. Section 3 describes our methodology for activation classification and Section 4 details our LLM fine tuning methodology for increasing the accuracy of activation classification. Section 5 discusses our results and final remarks are presented in Section 6.
2 Related Works
2.1 Previous Works Utilizing Activations
As seen in Tenney et al. (2019), BERT, a encoder-only LLM, is able to act as a high-performance feature extractor for text data through extracting activations [6]. With their procedure, they are able to evaluate layer-by-layer performance on a variety of standard natural language processing metrics. These results provide the theoretical foundation for our experiments: can decoder-only LLMs like LLaMA 2 7B be used as feature extractors, for the express purpose of classifying harmful prompts?
@@ 2.1 Previous Works Utilizing Activations
Templeton et al (2024) trained autoencoders on the Claude 3 LLM activation space in order to discover features of abstract concepts, including concepts such as "security vulnerabilities in code" and "bias" [9]. We use a simpler model to classify a spectrum of attack behaviors as specified by a potential user. This specification is provided through a example prompts, rather than abstract concepts.
2.2 Comparison to Previous Detection Methods
One way to detect attack prompts is by monitoring the perplexity measurement of any input prompt into an LLM. Perplexity is a measurement of fluency of the prompt, with high measurements indicating that a sequence of words is not likely to have been observed in the corpus the LLM was trained on. Attacks that score high in perplexity are known as token-level attacks, attacks generated through a reverse optimization of the embedding-space representation of words in the LLM. The following is an example of a successful token-level attack on Vicuna 7B [10]:
[…]
Jain et al. (2023) provides an example of a detection methodology that utilizes this measurement and achieves 100% detection rate on attack prompts of this type [11]. While this detection method is highly accurate for token-level attacks, it does not detect other types of attack prompt. Our defense aims to successfully detect all attack types.
The attack detection framework described in Pisano et al. (2023) utilizes a secondary LLM that monitors input prompts to the original LLM and verifies the corresponding text generation. The authors test this framework with a variety of attack styles and models. This framework achieves up to 95% detection rates of attacks with as low as a 7% false positive rate, depending on the type of model used as the secondary LLM [12]. While this framework has high detection accuracy, the computational cost incurred by using two LLMs can be prohibitively expensive. Our defense aims to achieve similarly accurate results without the use of a second LLM.
3 Methodology - Activation Classification
In this section, we define residual stream activations in a transformer model (3.1) and detail the steps we take to analyze and classify them into Jailbreak or Benign prompts (3.2). We also describe the malicious and benign prompt datasets we use (3.3) and the LLMs we use in the experiments (3.5). Code to reproduce this methodology is provided at the author’s repository  [^1] https://github.com/amelia-kawasaki/llm_activation_classification
InputTransformer Layer 1+Transformer Layer 2+Transformer Layer 3+Residual Stream
@@ 3 Methodology - Activation Classification
3.1 Residual Stream Activations
@@
3.2 Activation Analysis Methodology
@@
    Activation Vector Collection: For a prompt $P_{m}$ with $N$ tokens $t_{1},t_{2},\ldots,t_{N}$, where $m$ indexes the prompt from 1 to $M$ (the total number of prompts in a dataset), we collect activation vectors from each layer. For each token $t_{k}$ in prompt $P_{m}$, we obtain 32 activation vectors (one per layer), denoted as $v_{1}^{(t_{k})},v_{2}^{(t_{k})},\ldots,v_{32}^{(t_{k})}$.
@@
    $v_{1}^{(m)}=\text{average}(v_{1}^{(t_{1})},v_{1}^{(t_{2})},\ldots,v_{1}^{(t_{N})}),$
@@
    $v_{32}^{(m)}=\text{average}(v_{32}^{(t_{1})},v_{32}^{(t_{2})},\ldots,v_{32}^{(t_{N})}).$
    This averaging ensures that each set of activations for prompt $P_{m}$ has consistent dimensions regardless of the prompt’s length. Figure 2 provides an example of this process for the first two transformer layers in an LLM.
@@
    Classifier Training: We train a LightGBM classifier for each set of layer vectors to determine if the activations can be differentiated by class [5]. In this case, we train 32 classifiers, $\{C_{1},C_{2},\ldots,C_{32}\}$. Classifier $C_{i}$ uses set $V_{i}$ for training, where $V_{i}$ is the set of all $v_{i}^{(m)}$ vectors for every prompt $P_{m}$. The classifiers are configured with a maximum depth of 6 and 100 estimators, except for the WildJailBreak dataset in which we use a grid search to fit parameters for each classifier. We reserve 10% of the activation dataset for testing to verify the generalizability of the classifiers and to check for overfitting.
In order to provide a baseline for comparison, for some of our experiments we also collected activations before the prompt input is processed by the first transformer layer. At this point in the LLM, the prompt would have undergone a variety of pre-processing steps, including tokenization, embedding, and positional encoding. These activations are also used to train a classifier and the results are provided in Appendix B as the “layer 0” entry for each table.
3.3 Data Sources
For our analysis, we use 3 different datasets, each encompassing a different scope of attack prompts. Our intention is to simulate attacks that make up a representative selection of the spectrum of possible attack domains. The different types of attacks are: Broad, Domain-Specific, and Hyper-Specific. The Broad category consists of attacks that attempt to engage an LLM in a broad spectrum of undesirable behaviors. The Domain-Specific category consists of attacks pertaining to a single domain, simulating attacks on an LLM that is fine-tuned for a specific task. The Hyper-Specific category consists of attacks that aim at producing a specific undesirable string in an LLM’s response. Notably, we did not require that all attack prompts be successful. This is because an ideal LLM defense system should be able to detect on-going attacks on the model even if they are unsuccessful, so counter-measures can be taken. That said, we do test our classifier on an LLM that underwent fine-tuning to disengage safety training and on which all attacks were successful, further detailed in 3.5.
[…]
Figure 2: A subset of transformer layers of an LLM. For a given prompt $m$ with n number of tokens, an activation matrix is extracted after each transformer layer. This matrix is averaged across the tokens with the final averaged activation vector of size 1-by-size of hidden layer of the given LLM.
3.3.1 Broad Category
The attack dataset we use to encompass a broad range of attacks is the JailbreakV-28K dataset provided by Luo et al. (2024) available on HuggingFace [13]. We use the jailbreak_query column as the attacks. The dataset contains categories for each type of attack such as “Fraud”, “Animal Abuse”, and “Malware”, among others, which demonstrates the large variance of harmful topics for attack prompt generation. The class of benign prompts originates from Open-Orca created by Mukherjee et al. (2023) available on HuggingFace [14]. Additionally, we run a limited experiment which removes one of the attack types from the training set and reserve it for the testing set, in order to test if this detection procedure will generalize to unseen attacks. To do this, we withheld attacks with “Persuade” listed in the format column. These type of attacks attempt to reason or threaten the LLM into generating harmful output.
Although both of the original datasets have over 100,000 prompts each, we use a subset of the dataset for analysis due to time and computation constraints. The subset for our analysis contains 50,000 prompts split evenly between attack and benign prompts. We also reserve a smaller subset of the prompts for our fine tuning procedure: another 2,000 prompts split evenly between attack and benign prompts.
3.3.2 Domain-Specific Category
Table 1: Prompt Length Restriction by Dataset
@@ 3.3.2 Domain-Specific Category
Domain-Specific Attack [10-845] [50-170] 3057
Benign [9-438] [50-170] 3057
Hyper-Specific: Attack [63-606,391] [500-1500] 5358
HackAPrompt 2023 Benign [22-35,871] [500-1500] 5358
3.3.3 Hyper-Specific Category
We have two datasets for this category that target classification of specific attacks. The first attack dataset for this analysis originates from submissions to HackAPrompt 2023, an online competition to create successful attack prompts [16]. The participants in this competition created prompts that attempted to make an LLM output the string:
@@ 3.3.3 Hyper-Specific Category
which is the goal output string that indicates that an LLM has been compromised. The attack dataset was then made open source on HuggingFace, a popular website for sharing models and datasets [17]. We use a subset of the original dataset, totaling 25,000 attacks. For the benign prompts, we use the same subset as the Broad dataset: 25,000 prompts from the Open-Orca dataset. We also use another smaller subset of this dataset to create the Hyper-Specific data for our fine tuning procedure, another 2,000 prompts split evenly between attack and benign prompts.
The second dataset for this analysis is the WildJailbreak set provided by Jiang et al. (2024) available on HuggingFace [18]. This dataset provides benign and attack prompts that all utilize a role-playing style in order to make a specific request to the LLM. While the other datasets we test on have easily distinguishable classes of benign and attack prompts, this dataset’s classes are similar in structure and topic. We included this dataset in order to test our methodology’s ability to classify nuanced prompts. We use a subset of the original dataset, totaling 3,000 attacks split between classes. For each classifier trained on the activations of this data, we use a grid search to optimize parameters in order to maximize the classifier’s ability to differentiate between classes. We did not include this dataset in our fine-tuning methodology.
3.4 Dataset Modification
We additionally perform prompt length range restriction on some of the datasets. The prompt classes in the Hyper-Specific and Broad datasets have different distributions of prompt length. In order to avoid prompt length becoming a determining feature of the classifier, we make modified versions of these datasets with the range of the prompt lengths restricted. The new versions of these datasets are subsets of the original datasets that only contain prompts that match the requirements for length. The new dataset is then class balanced by removing a random selection of prompts from the larger class. We perform our analysis with the original and range-restricted datasets and report results from each. Table 1 provides information on the transformations to each dataset. We determined that the prompt length range for the Domain-Specific dataset was similar enough between prompt classes and did not need any modification. The Wildjailbreak dataset is not included in this set of modifications.
3.5 Models
@@
4 Methodology - Fine Tuning for Improved Classification
With the intention of improving attack prompt detection accuracy, we fine tune LLaMA 2 7B using Low Rank Approximation of LLMs (LoRA) [23]. This lightweight and parameter efficient training procedure allows us to fine tune the 7B model on one GPU. In total, we fine tune three versions of LLaMA 2 7B, once for each of the dataset types we detailed in 3.3. For each model, we use a hold-out set of prompts solely used for fine tuning (2,000 for Broad, 1,000 for Domain-Specific, 2,000 for Hyper-Specific, each split evenly across classes). Taking inspiration from Raman and Maini (2023), we then apply a format to each prompt in order to force the model to evaluate each prompt in the context of detecting prompt injections [24]. We then provide responses to each prompt for the LLM to emulate. The LLM is then trained on the combination of augmented prompt and response to complete the LoRA fine tuning procedure. To make the augmented prompt, we started by prepending the string:
@@ 4 Methodology - Fine Tuning for Improved Classification
to the end of the prompt. This frames the original prompt in an appropriate context and asks the LLM to only respond with “yes” or “no”. We then provide a corresponding response that the LLM should train to emulate. The response is “yes” if the original prompt belongs to the attack prompt class or “no” if the original prompt belongs to the benign prompt class. After fine tuning the three versions of LLaMA 2 7B, we repeat the activation classification procedure we detailed in 3.2 and compare the results to the results utilizing non-fine tuned LLMs.
5 Results
Table 2: Broad Attack Results - Averaged Over Layers
Model Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
LLaMA 2 7B 0.999$\pm$0.0001 0.001 0.001
LLaMA 2 13B 0.999$\pm$0.0005 0.001 0.001
TinyLlama 1.1B 0.999$\pm$0.0004 0.001 0.001
Mistral 7B 0.999$\pm$0.0001 0.001 0.001
Vicuna 7B 0.999$\pm$0.0001 0.001 0.001
Wizard Vicuna 7B Uncensored 0.999$\pm$0.0005 0.001 0.001